Two small copper spheres of mass 0.0180 kg are 50 cm apart. a) How many electrons does each sphere contain? (copper atomic number is 29 and the molar mass is 63.546 g/mol). b) How many electrons would have to be removed (from one sphere) and added (to the other) such that an attractive force of 3x10^2 N is generated? Treat the spheres as point charges.
1
Expert's answer
2021-10-13T15:42:33-0400

a)


"N=(6.02\cdot10^{23})\frac{0.018}{63.546}=1.705\cdot10^{20}"

b)


"300=(9\cdot10^9)\frac{(1.6\cdot10^{-19})^2n^2}{0.5^2}\\n=5.705\cdot10^{14}"